Ill ule pri 01 0.010 VCICIO4)3 4. A 0.0678 M solution of an acid is 0.55% dissociated. Calculate the pka for this a
Answer #1

HA - 0.06 77 H+ + A- - - 1 x x E 0.0678.7 0.55 0.0678 loo - 3.73610-4 Ka= [H] [A-] THAI 22 0.0 678 - X = (3.73*10-4) 2 0.0 67